OpenStudy (anonymous):

BASIC VECTOR QUESTION: The pilot of an airplane that flies at 800 km/h wishes to travel to a city 800 km due east. There is a 80 km/h wind from the north east. What should the plane's heading be?

OpenStudy (kainui):

Draw a picture, and make your best guess.

OpenStudy (anonymous):

i did, five me a sec.

OpenStudy (kainui):

Sure! =)

OpenStudy (anonymous):

 i assume that resultant velocity will be in the east since the pilot "WISHES" to fly in east.. then velocity of the wind is "FROM" north east.. so i calculated for \( \alpha\) using sine law then i calculated the \(\theta = 180-135-4.05=40.9 \), which is the plane's direction, but it's not the right answer then i have another solution which i get the right answer but it doesn't make sense:/ in my second solution i assume that 800km/h is the velocity of the plane without any wind affecting it